OpenStudy (anonymous):

Solve the equation. 78 = -2(m + 3) + m A. –28 B. –42 C. –72 D. –84

OpenStudy (anonymous):

78=-2m-6+m=-m-6 m=-78-6=-84

OpenStudy (anonymous):

thanks